Find the measure of an angle that is twice as large as its complement.
A supplement angle is four times as large as the angle. Find the measure of the angle.
The measure of a complement of an angle is three times more than twice the measure of the angle. Find the measure of the angle and its complement.

Answer by BurningFlame_marc(77)

Hello!
I'll just answer you're questions ok?
1.Find the measure of an angle that is twice as large as its complement.
Representation: let x be the measure of the angle's complement,and 2x be the measure of the angle.
Equation: 2x+x=90
Solution:
3x=90
x=30
Final Answer: 30
2.A supplement angle is four times as large as the angle. Find the measure of the angle.
Representation: let x be the measure of the angle,and 4x be the measure of the angle's supplement.
Equation: 4x+x=180
Solution:
5x=180
x=36
Final Answer: 36
3. The measure of a complement of an angle is three times more than twice the measure of the angle. Find the measure of the angle and its complement.
Representation: let x be the measure of the angle,and 3(2x)+2x be the measure of the angle's complement.
Equation: 3(2x)+2x+x=90
Solution:
9x=90
x=10
Final Answer: the angle is 10,the complement is 80.

*If you have any questions feel free to ask me. Just email me at BurningFlame_marc@yahoo.com. Thanks! Have a nice day!